# Analytically Continuing the Randomized Measurement Toolbox

### v1

@@ Appendix A Stabilized Analytic Continuation
In this appendix, we review the key features of the SAC method outlined in references [38, 39, 40, 41, 42]. Additional details can be found in the original references.
The basic setup of the problem is as described in the main text. $F(w)$ is an unknown analytic function that is holomorphic in the interior of the unit disk. It assumes real values on the real axis: $F(w_{i})=a_{i}\in\mathbb{R}$ for $z_{i}\in(-1,1)$ and i=2,3,,N+1i=2,3,...,N+1, where $N$ denotes the number of data points ($N=k_{\text{max}}-1$ in the main text). Our aim is to fit for the value of the function at a specific point on the unit circle, which we take to be w=1w=-1 without loss of generality. SAC was originally developed by Ciulli and Spearman [38, 39, 40, 41, 42] to diagnose resonances of the Green’s function which lie outside, but close to the unit disk, using a finite set of noisy data points provided within the interior of the disk. The presence of nearby poles results in a significant amount of structure in the function on the unit circle, which can be measured by a suitable choice of norm. Minimizing this norm over all consistent analytic continuations then yields the minimal amount of structure that is forced upon the function by the data itself.
@@
As illustrated in [38, 39, 40, 41, 42], knowing $x_{,r}(\theta)$ on the unit circle allows one to reconstruct the function $X(w)$ in the interior of the unit disk upto an overall constant through the formula
@@
at the data points $\{w_{i}\}$. There is an infinite family of analytic functions $D^{\prime}_{\alpha}(w)$ that satisfy the condition. From this infinite family, we need to select the analytic continuation that has the smallest norm. As illustrated in [38, 39, 40, 41, 42], this non-linear minimization problem can be recast into a linear matrix optimization problem, and the minimal norm (squared) $\delta_{\min}^{2}$ of $D^{\prime}_{\alpha}(w)$ is given by
@@ Appendix B Stabilized Analytic Continuation with Errors
Due to errors in the data points, it now no longer makes sense to choose the subtraction point to correspond to one of the data points. To incorporate errors, the subtraction point can instead be chosen as the guess function’s value at $w_{2}$ [39], or as an additional variational parameter [42]. Here we opt for the latter and take $w_{0}$, an arbitrary point on the real interval (1,1)(-1,1), as our subtraction point. Given any data point $\mathbf{y}$ within this ellipse $\chi^{2}\leq\chi^{2}_{0}$, the minimal (squared) norm $\delta^{2}_{\min}$ is given by

### v2

@@ Appendix A Stabilized Analytic Continuation
In this appendix, we review the key features of the SAC method outlined in references [CS1, CS2, CS3, CS4, CS5]. Additional details can be found in the original references.
The basic setup of the problem is as described in the main text. $F(w)$ is an unknown analytic function that is holomorphic in the interior of the unit disk. It assumes real values on the real axis: $F(w_{i})=a_{i}\in\mathbb{R}$ for $z_{i}\in(-1,1)$ and i=2,3,,N+1i=2,3,...,N+1, where $N$ denotes the number of data points ($N=k_{\text{max}}-1$ in the main text). Our aim is to fit for the value of the function at a specific point on the unit circle, which we take to be w=1w=-1 without loss of generality. SAC was originally developed by Ciulli and Spearman [CS1, CS2, CS3, CS4, CS5] to diagnose resonances of the Green’s function which lie outside, but close to the unit disk, using a finite set of noisy data points provided within the interior of the disk. The presence of nearby poles results in a significant amount of structure in the function on the unit circle, which can be measured by a suitable choice of norm. Minimizing this norm over all consistent analytic continuations then yields the minimal amount of structure that is forced upon the function by the data itself.
@@
As illustrated in [CS1, CS2, CS3, CS4, CS5], knowing $x_{,r}(\theta)$ on the unit circle allows one to reconstruct the function $X(w)$ in the interior of the unit disk upto an overall constant through the formula
@@
at the data points $\{w_{i}\}$. There is an infinite family of analytic functions $D^{\prime}_{\alpha}(w)$ that satisfy the condition. From this infinite family, we need to select the analytic continuation that has the smallest norm. As illustrated in [CS1, CS2, CS3, CS4, CS5], this non-linear minimization problem can be recast into a linear matrix optimization problem, and the minimal norm (squared) $\delta_{\min}^{2}$ of $D^{\prime}_{\alpha}(w)$ is given by
@@ Appendix B Stabilized Analytic Continuation with Errors
Due to errors in the data points, it now no longer makes sense to choose the subtraction point to correspond to one of the data points. To incorporate errors, the subtraction point can instead be chosen as the guess function’s value at $w_{2}$ [CS2], or as an additional variational parameter [CS5]. Here we opt for the latter and take $w_{0}$, an arbitrary point on the real interval (1,1)(-1,1), as our subtraction point. Given any data point $\mathbf{y}$ within this ellipse $\chi^{2}\leq\chi^{2}_{0}$, the minimal (squared) norm $\delta^{2}_{\min}$ is given by